OpenStudy (anonymous):

In a 1224 km car race, Peter drove at a speed of 180 km/h while David's speed was 10 km/h less. How much sooner did Peter finish the race than David?

OpenStudy (anonymous):

d=vt t=d/v Time for Peter = 1224/180 Time for David = 1224/170 Difference in time, in hours, between David and Peter: 1224/170-1224/180

OpenStudy (anonymous):

7.2-6.8= .4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

He finished .4 hours ahead of David, or 24 minutes

OpenStudy (anonymous):

thanks cause i didn't really get it because i did it the other way around.

OpenStudy (anonymous):

You could've gotten the negative 0.4, it just tells you he finished ahead of David